OpenStudy (anonymous):

sovle the system of equations 3x-9y=72 2x-6y=-14

OpenStudy (anonymous):

mutliply the first equation by 2/3 to get 2x-6y=48 2x-6y=-14 look at the left sides... the slopes of the lines are the same, but (look at the right side of the equations) the y-intercepts are different, ie, the lines are parallel => no solution to this system

OpenStudy (anonymous):

x=-93/2 ==> y =-31/2

OpenStudy (anonymous):

the first equation in slope intercept form is\[y=\frac{1}{3}x-8\]the second line is\[y=\frac{1}{3}x+\frac{7}{3}\]you can see that the slope are equal but they intercect the y-axis at different locations. Parallel lines do not intersect and that is the geometric interpretation of solving this kind of system.

OpenStudy (anonymous):

that is the point or points of intersection is(are) the solution

OpenStudy (anonymous):

here is the graph of the two lines